# Connectivity Structure and Dynamics of Nonlinear Recurrent Neural Networks

### v2

@@ Appendix A Glossary of terms
Number of neurons ($N$)
Total number of neurons in the network.
Number of modes ($M$)
Number of components in the random-mode model.
(Effective) coupling strength ($g$, $g_{\text{eff}}$)
Standard deviation of the couplings times $\sqrt{N}$ in the i.i.d. connectivity model, random-mode model (Eq. 37), or variants thereof (e.g., see just below Eq. 46a and Eq. 46b).
Random-mode model
Network connectivity model: $\bm{J}=\sum_{a=1}^{M}D_{a}\bm{\ell}_{a}\bm{r}_{a}^{T}$ (Eq. 2).
Connectivity component ($\bm{\ell}_{a}\bm{r}_{a}^{T}$)
Outer product of output and right modes.
Right mode ($\bm{r}_{a}$)
Neuronal pattern in the $a$-th component onto which activity patterns are projected.
Left mode ($\bm{\ell}_{a}$)
Neuronal pattern in the $a$-th component along which the projected activity pattern is expanded.
Component strength ($D_{a}$)
Scaling factor for the $a$-th component.
Ratio of modes to neurons ($\alpha$)
Given by $\alpha=M/N$ (Eq. 7).
Participation ratio of component strengths ($\text{PR}^{D}$)
Given by $\text{PR}^{D}=r_{2}^{2}/r_{4}$ (Eq. 11) where $r_{n}=\langle D^{n}\rangle_{D}$ (Eq. 10).
Effective rank ($\alpha\text{PR}^{D}$)
Connectivity dimensionality measure (Eq. 12).
Two-point function
Correlation function given by $C^{\phi}(t_{1},t_{2})=\frac{1}{N}\sum_{i=1}^{N}\langle\phi_{i}(t_{1})\phi_{i}(t_{2})\rangle_{t}$ (Eq. 17).
Four-point function
Higher-order correlation function given by $\Psi^{\phi}(\tau_{1},\tau_{2})=\frac{1}{N}\sum_{i,j=1}^{N}C^{\phi}_{ij}(\tau_{1})C^{\phi}_{ij}(\tau_{2})$ (Eq. 24). See also time-by-time definition (Eq. 25).
@@ Appendix B Preprocessing of fly hemibrain connectome
GABA -1 (Inhibitory)
Acetylcholine +1 (Excitatory)
Glutamate -1 (Inhibitory)
@@
Table 1: Mapping of neurotransmitters to synaptic signs
@@
  • the sparsity pattern (which connections are zero vs. nonzero);
  • the sign of each connection (excitatory vs. inhibitory);
@@
We normalize the coupling matrix using an iterative scaling method analogous to the Sinkhorn–Knopp algorithm. First, for each row, we rescale the positive elements by a row-specific factor so that their L2-norm equals $1/\sqrt{2}$. We then rescale the negative elements, using a separate factor, so that their L2-norm also equals $1/\sqrt{2}$. This yields rows with a total L2-norm of one and balanced excitatory and inhibitory contributions. A small number of neurons with only excitatory or only inhibitory inputs retain an L2-norm of $1/\sqrt{2}$.
@@ Appendix C Free probability calculation
which is the product of two independent Wishart matrices:
@@
In a 2D electrostatics analogy, the real and imaginary parts of $G(z)$ correspond to the Cartesian components of the electric field generated by a line charge density $\rho(\lambda)$ on the real axis. The inversion formula is then analogous to Gauss’s law, relating the discontinuity of the normal field component to the local charge density.
The moment generating function and S-transform are defined by
@@
where $M^{-1}(z)$ is the functional inverse of $M(z)$. For a Marchenko–Pastur distributed Wishart matrix $\bm{W}$ with aspect ratio $\alpha$,
@@
A key fact in free probability is that the S-transforms of free variables multiply. Thus, for $\bm{W}_{1}\bm{W}_{2}$, the S-transform is
@@
From the definition of the S-transform, the inverse moment generating function is
@@
$z=\frac{(1+M(z))(1+\alpha M(z))^{2}}{M(z)}.$ (63)
Using M(z)=zG(z)1M(z)=zG(z)-1, we obtain a cubic equation for $G(z)$:
@@
In principle, this can be solved using the cubic formula, but the explicit form is cumbersome. For the support boundaries of $\rho(\lambda)$, we only need the points where $\mathrm{Im}G(z)$ changes from zero to nonzero. This occurs when the discriminant of the cubic vanishes. The discriminant is
@@
The singular values correspond to the square roots of the eigenvalues, given by Eq. 14.
Figure 8: Distribution of singular values in the random-mode model. Histograms show singular values of $\bm{L}\bm{D}\bm{R}^{T}$ with $D_{a}=1$ for various $\alpha$. For $\alpha\ll 1$, the singular values concentrate at one. Vertical lines: prediction of Eq. 14. $N=5000$.
@@ Appendix D Alternative measures of effective dimensionality
Alternative measures of effective dimensionality, which could be applied to both connectivity and activity in place of the participation ratio, have been proposed in the literature. Roy and Vetterli [88] define effective rank as the exponential of the entropy of a categorical probability distribution obtained by normalizing the eigenvalues values to sum to one, i.e., $p_{i}=\lambda_{i}/\sum_{j=1}^{N}\lambda_{j}$. Written in terms of this distribution the participation ratio is $\text{PR}=\left(\sum_{i=1}^{N}p_{i}^{2}\right)^{-1}$, while the exponentiated entropy is $\exp(H)=\exp\left(-\sum_{i=1}^{N}p_{i}\log p_{i}\right)$. By Jensen’s inequality, $\text{PR}\leq\exp(H)$. Exact equality holds in certain cases, including when exactly $K\leq N$ eigenvalues are equal and the rest are zero, yielding $\text{PR}=\exp(H)=K$. Furthermore, for smoothly decaying spectra as discussed above (for which $p_{i}=f(i/Nw)/\sum_{j=1}^{N}f(j/Nw)$, with $w\ll 1$), both measures have $\text{PR}=c_{1}\>Nw$ and $\exp(H)=c_{2}\>Nw$, where $c_{1}$ and $c_{2}$ are $N$ and $w$-independent constants depending on the specific form of $f$. Thus, while the measures differ in detail, they provide similar characterizations of effective dimensionality. This entropy-based definition has been used in neuroscience applications, for example by [89] to analyze transient amplification in neural networks. Both measures capture the number of dominant modes in the system. Thibeault et al. [19] call (a metric closely related to) the the participation ratio-based metric the “srank” (the difference being whether the maximum eigenvalue or sum over all eigenvalues is used as a normalizing factor) and call the entropy-like metric the “erank.”
@@
We now review how to compute these two- and four-point functions in a classic network model with i.i.d. couplings $\bm{J}$. The first- and second-order coupling statistics are:
@@
The magnitude of a typical coupling is $g/\sqrt{N}$. We refer to $g$ as the coupling strength. For the canonical dynamics $T[x](t)=(1+\partial_{t})x(t)$, the network is quiescent for $g<1$ and chaotic for $g>1$, with a sharp phase transition as $N\to\infty$. In this paper we assume that the network is non-quiescent.
The two-point function $C^{a}_{\star}(t_{1},t_{2})$ can be computed through a single-site picture that describes the dynamics of a typical neuron embedded in the rest of the network, with pre-activation $x(t)$ and activation $\phi(t)$. The single-site dynamics are given by
@@
where $\left\langle\cdots\right\rangle_{\star}$ denotes an average within this single-site process, i.e., with respect to the Gaussian distribution of $\eta^{x}(t)$. Once $C^{\phi}_{\star}(t_{1},t_{2})$ has been determined, $C^{x}_{\star}(t_{1},t_{2})$ follows easily. This single-site problem can be derived through either a single-site cavity calculation (see [90]) or a saddle-point condition in a path integral (see [54, 55]). For the i.i.d. couplings considered here, there is a simpler heuristic derivation: in the neuronal input $\sum_{j=1}^{N}J_{ij}\phi_{j}(t)$, the correlations between the couplings $J_{ij}$ and dynamic variables $\phi_{j}(t)$ can be safely neglected to leading order in $1/N$, yielding both Gaussianity of $\eta^{x}(t)$ by the central limit theorem and the second-order statistics of Eq. 70.
Clark et al. [30] first computed $\Psi_{\star}^{a}(\bm{\tau})$ using a dynamic, two-site version of the cavity method, based on the neuron-by-neuron definition (Eq. 24). This method finds $\psi^{a}_{\star}(\bm{\tau})$, the off-diagonal contribution to $\Psi^{a}_{\star}(\bm{\tau})$, noting that the on-diagonal contribution is simply $C^{a}_{\star}(\tau_{1})C^{a}_{\star}(\tau_{2})$. A cavity is first created by removing two neurons from the network and allowing the rest of the network, the reservoir, to generate dynamic activity. The cavity neurons are then introduced, and their effect on the reservoir is treated perturbatively. This yields a pair of coupled mean-field equations for the cavity units, generalizing the single-site picture discussed above to a two-site picture. Finally, self-consistency conditions are constructed by recognizing that the cavity pair is statistically equivalent to any reservoir pair. This calculation results in expressions for the four-point function in Fourier space, given by
@@
for the pre-activations, where $\bm{\omega}=(\omega_{1},\omega_{2})$ and $C^{x\phi}_{\star}(\omega)$ is a cross-covariance between the pre-activation and activation. Note that if the joint distribution of pre-activations were Gaussian, $\Psi^{x}_{\star}(\bm{\tau})$ and $\Psi^{\phi}_{\star}(\bm{\tau})$ would differ only by a proportionality constant due to Price’s theorem [91]. The more complex relationship observed here reflects the non-Gaussian joint statistics across different neurons, which is relevant because the network is nonlinear.
@@
Figure 9: Relationship between principal component timescales and two- and four-point correlation functions. (a) Example activity traces of principal components $p^{\phi}_{k}(t)$. (b) Empirical autocovariance functions for each principal component, weighted by their eigenvalues. (c) Same as (b) but weighted by the squared eigenvalue. (d) Averages over curves in (b) and (c), with comparisons to $C^{\phi}_{\star}(\tau)$ and $\Psi^{\phi}_{\star}(\tau,0)$. Dots: simulation results. Lines: theoretical predictions. All panels describe the same network of $N=4000$ neurons with i.i.d. couplings with variance $g^{2}/N$ with coupling strength $g=6$.
@@
    Formulate the field theory. Begin with the path integral $Z[\bm{J}]$ and average over the connectivity disorder. Introduce auxiliary fields representing two-point functions (such as $C^{\phi}(t_{1},t_{2})$) along with their conjugate partners. Use integral representations of delta functions to enforce field definitions, factorizing the exponential over extensive dimensions. This yields a statistical field theory with action $N$ times an order-one function of the auxiliary fields.
@@
    Solve the saddle-point equations. Find the saddle point by setting action derivatives to zero. This produces self-consistent equations for the two-point functions, completing the standard dynamical mean-field theory (DMFT) analysis.
@@
    Apply temporal separation limits. Impose the temporal limits (Eq. 34), making the Hessian blocks translation-invariant functions of time differences $(\tau_{1},\tau_{2})$. Transform to Fourier space to obtain a low-dimensional, frequency-dependent Hessian whose elements are functions of $(\omega_{1},\omega_{2})$.
@@
Here we consider i.i.d. $\bm{J}$ with mean zero and variance $g^{2}/N$. After performing the Gaussian integration over $\bm{J}$, we introduce an auxiliary field $C^{\phi}(t_{1},t_{2})$, defined by Eq. 17, and its conjugate $\hat{C}^{\phi}(t_{1},t_{2})$ to factorize the action across the neuron index $i$. This leads to a partition function for a statistical field theory involving $C^{\phi}(t_{1},t_{2})$ and $\hat{C}^{\phi}(t_{1},t_{2})$:
@@ F.1 Two-point functions
where $\left\langle\cdots\right\rangle_{\star}$ denotes an average within the dynamic process described by $W[C_{\star}^{\phi},0]$. This recovers the same single-site process (Eq. 69, Eq. 70) and self-consistency condition (Eq. 71) described in Sec. E.

### v3

@@ Appendix A Glossary of terms
Number of neurons ($N$): Total number of neurons in the network.
Number of modes ($M$): Number of components in the random-mode model.
(Effective) coupling strength ($g$, $g_{\text{eff}}$): Standard deviation of the couplings times $\sqrt{N}$ in the i.i.d. connectivity model, random-mode model [Eq. (38)], or variants thereof [e.g., see just below Eqs. (47a) and (47b)].
Random-mode model: Network connectivity model: $\bm{J}=\sum_{a=1}^{M}D_{a}\bm{\ell}_{a}\bm{r}_{a}^{T}$ [Eq. (2)].
Connectivity component ($\bm{\ell}_{a}\bm{r}_{a}^{T}$): Outer product of output and right modes.
Right mode ($\bm{r}_{a}$): Neuronal pattern in the $a$th component onto which activity patterns are projected.
Left mode ($\bm{\ell}_{a}$): Neuronal pattern in the $a$th component along which the projected activity pattern is expanded.
Component strength ($D_{a}$): Scaling factor for the $a$th component.
Ratio of modes to neurons ($\alpha$): Given by $\alpha=M/N$ [Eq. (7)].
Participation ratio of component strengths ($\text{PR}^{D}$): Given by $\text{PR}^{D}=r_{2}^{2}/r_{4}$ [Eq. (11)] where $r_{n}=\langle D^{n}\rangle_{D}$ [Eq. (10)].
Effective rank ($\alpha\text{PR}^{D}$): Connectivity dimensionality measure [Eq. (12)].
Two-point function: Correlation function given by $C^{\phi}(t_{1},t_{2})=(1/N)\sum_{i=1}^{N}\phi_{i}(t_{1})\phi_{i}(t_{2})$ [Eq. (17)].
Four-point function: Higher-order correlation function given by $\Psi^{\phi}(\tau_{1},\tau_{2})=(1/N)\sum_{i,j=1}^{N}C^{\phi}_{ij}(\tau_{1})C^{\phi}_{ij}(\tau_{2})$ [Eq. (24)]. See also time-by-time definition [Eq. (25)].
@@ Appendix B Preprocessing of fly hemibrain connectome
GABA 1-1 (Inhibitory)
Acetylcholine +1+1 (Excitatory)
Glutamate 1-1 (Inhibitory)
@@
Table 1: Mapping of neurotransmitters to synaptic signs.
@@
  1. (i)
    the sparsity pattern (for which the connections are zero versus nonzero);
  2. (ii)
    the sign of each connection (excitatory versus inhibitory);
  3. (iii)
@@
We normalize the coupling matrix using an iterative scaling method analogous to the Sinkhorn-Knopp algorithm. First, for each row, we rescale the positive elements by a row-specific factor so that their $L2$ norm equals $1/\sqrt{2}$. We then rescale the negative elements, using a separate factor, so that their $L2$ norm also equals $1/\sqrt{2}$. This yields rows with a total $L2$ norm of one and balanced excitatory and inhibitory contributions. A small number of neurons with only excitatory or only inhibitory inputs retain an $L2$ norm of $1/\sqrt{2}$.
@@ Appendix C Free probability calculation
which is the product of two independent Wishart matrices,
@@
In a two-dimensional electrostatics analogy, the real and imaginary parts of $G(z)$ correspond to the Cartesian components of the electric field generated by a line charge density $\rho(\lambda)$ on the real axis. The inversion formula is then analogous to Gauss’s law, relating the discontinuity of the normal field component to the local charge density.
The moment-generating function and $S$ transform are defined by
@@
where $M^{-1}(z)$ is the functional inverse of $M(z)$. For a Marchenko-Pastur distributed Wishart matrix $\bm{W}$ with aspect ratio $\alpha$,
@@
A key fact in free probability is that the $S$ transforms of free variables multiply. Thus, for $\bm{W}_{1}\bm{W}_{2}$, the $S$ transform is
@@
From the definition of the $S$ transform, the inverse moment-generating function is
@@
$z=\frac{[1+M(z)][1+\alpha M(z)]^{2}}{M(z)}.$ (64)
Using M(z)=zG(z)1M(z)=zG(z)-1, we obtain a cubic equation for $G(z)$,
@@
In principle, this can be solved using the cubic formula, but the explicit form is cumbersome. For the support boundaries of $\rho(\lambda)$, we need only the points where $\mathrm{Im}G(z)$ changes from zero to nonzero. This occurs when the discriminant of the cubic vanishes. The discriminant is
@@
The singular values correspond to the square roots of the eigenvalues, given by Eq. (14). We confirm this formula numerically in Fig. 8.
Figure 8: Distribution of singular values in the random-mode model. Histograms show singular values of $\bm{L}\bm{D}\bm{R}^{T}$ with $D_{a}=1$ for various $\alpha$. For $\alpha\ll 1$, the singular values concentrate at 1. Vertical lines, prediction of Eq. (14). $N=5000$.
@@ Appendix D Alternative measures of effective dimensionality
Alternative measures of effective dimensionality, which could be applied to both connectivity and activity in place of the participation ratio, have been proposed in the literature. Roy and Vetterli [95] define effective rank as the exponential of the entropy of a categorical probability distribution obtained by normalizing the eigenvalues to sum to 1, i.e., $p_{i}=\lambda_{i}/\sum_{j=1}^{N}\lambda_{j}$. Written in terms of this distribution, the participation ratio is $\text{PR}=\left(\sum_{i=1}^{N}p_{i}^{2}\right)^{-1}$, while the exponentiated entropy is $\exp(H)=\exp\left(-\sum_{i=1}^{N}p_{i}\log p_{i}\right)$. By Jensen’s inequality, $\text{PR}\leq\exp(H)$. Exact equality holds in certain cases, including when exactly $K\leq N$ eigenvalues are equal and the rest are zero, yielding $\text{PR}=\exp(H)=K$. Furthermore, for smoothly decaying spectra as discussed above [for which $p_{i}=f(i/Nw)/\sum_{j=1}^{N}f(j/Nw)$, with $w\ll 1$], both measures have $\text{PR}=c_{1}\>Nw$ and $\exp(H)=c_{2}\>Nw$, where $c_{1}$ and $c_{2}$ are $N$- and $w$-independent constants depending on the specific form of $f$. Thus, while the measures differ in detail, they provide similar characterizations of effective dimensionality. This entropy-based definition has been used in neuroscience applications, for example, in Ref. [96] to analyze transient amplification in neural networks. Both measures capture the number of dominant modes in the system. Thibeault et al. (Ref. [23]) call (a metric closely related to) the the participation ratio-based metric the “srank” (the difference being whether the maximum eigenvalue or sum over all eigenvalues is used as a normalizing factor) and call the entropylike metric the “erank.”
@@
We now review how to compute these two- and four-point functions in a classic network model with i.i.d. couplings $\bm{J}$. The first- and second-order coupling statistics are
@@
The magnitude of a typical coupling is $g/\sqrt{N}$. We refer to $g$ as the coupling strength. For the canonical dynamics $T[x](t)=(1+\partial_{t})x(t)$, the network is quiescent for $g<1$ and chaotic for $g>1$, with a sharp phase transition as $N\to\infty$. In this paper we assume that the network is nonquiescent.
The two-point function $C^{a}_{\star}(t_{1},t_{2})$ can be computed through a single-site picture that describes the dynamics of a typical neuron embedded in the rest of the network, with preactivation $x(t)$ and activation $\phi(t)$. The single-site dynamics are given by
@@
where $\left\langle\cdots\right\rangle_{\star}$ denotes an average within this single-site process, i.e., with respect to the Gaussian distribution of $\eta^{x}(t)$. Once $C^{\phi}_{\star}(t_{1},t_{2})$ has been determined, $C^{x}_{\star}(t_{1},t_{2})$ follows easily. This single-site problem can be derived through either a single-site cavity calculation (see Ref. [97]) or a saddle-point condition in a path integral (see Refs. [60, 61]). For the i.i.d. couplings considered here, there is a simpler heuristic derivation: In the neuronal input $\sum_{j=1}^{N}J_{ij}\phi_{j}(t)$, the correlations between the couplings $J_{ij}$ and dynamic variables $\phi_{j}(t)$ can be safely neglected to leading order in $1/N$, yielding both Gaussianity of $\eta^{x}(t)$ by the central limit theorem and the second-order statistics of Eq. (71).
Clark et al. [39] first computed $\Psi_{\star}^{a}(\bm{\tau})$ using a dynamic, two-site version of the cavity method, based on the neuron-by-neuron definition [Eq. (24)]. This method finds $\psi^{a}_{\star}(\bm{\tau})$, the off diagonal contribution to $\Psi^{a}_{\star}(\bm{\tau})$, noting that the on-diagonal contribution is simply $C^{a}_{\star}(\tau_{1})C^{a}_{\star}(\tau_{2})$. A cavity is first created by removing two neurons from the network and allowing the rest of the network, the reservoir, to generate dynamic activity. The cavity neurons are then introduced, and their effect on the reservoir is treated perturbatively. This yields a pair of coupled mean-field equations for the cavity units, generalizing the single-site picture discussed above to a two-site picture. Finally, self-consistency conditions are constructed by recognizing that the cavity pair is statistically equivalent to any reservoir pair. This calculation results in expressions for the four-point function in Fourier space, given by
@@
for the preactivations, where $\bm{\omega}=(\omega_{1},\omega_{2})$ and $C^{x\phi}_{\star}(\omega)$ is a cross-covariance between the preactivation and activation. Note that, if the joint distribution of preactivations were Gaussian, $\Psi^{x}_{\star}(\bm{\tau})$ and $\Psi^{\phi}_{\star}(\bm{\tau})$ would differ only by a proportionality constant due to Price’s theorem [98]. The more complex relationship observed here reflects the non-Gaussian joint statistics across different neurons, which is relevant because the network is nonlinear.
@@
Figure 9: Relationship between principal component timescales and two- and four-point correlation functions. (a) Example activity traces of principal components $p^{\phi}_{k}(t)$. (b) Empirical correlation functions for each principal component (PC), weighted by their eigenvalues. (c) Same as (b) but weighted by the squared eigenvalue. (d) Averages over curves in (b) and (c), with comparisons to $C^{\phi}_{\star}(\tau)$ and $\Psi^{\phi}_{\star}(\tau,0)$. Dots, simulation results. Lines, theoretical predictions. All panels describe the same network of $N=4000$ neurons with i.i.d. couplings with variance $g^{2}/N$ with coupling strength $g=6$.
@@
    Formulate the field theory. Begin with the path integral $Z[\bm{J}]$ and average over the connectivity disorder. Introduce auxiliary fields representing two-point functions [such as $C^{\phi}(t_{1},t_{2})$] along with their conjugate partners. Use integral representations of $\delta$ functions to enforce field definitions, factorizing the exponential over extensive dimensions. This yields a statistical field theory with action $N$ times an order-one function of the auxiliary fields.
@@
    Solve the saddle-point equations. Find the saddle point by setting action derivatives to zero. This produces self-consistent equations for the two-point functions, completing the standard DMFT analysis.
@@
    Apply temporal separation limits. Impose the temporal limits [Eq. (35)], making the Hessian blocks translation-invariant functions of time differences $(\tau_{1},\tau_{2})$. Transform to Fourier space to obtain a low dimensional, frequency-dependent Hessian whose elements are functions of $(\omega_{1},\omega_{2})$.
@@
Here we consider i.i.d. $\bm{J}$ with mean zero and variance $g^{2}/N$. After performing the Gaussian integration over $\bm{J}$, we introduce an auxiliary field $C^{\phi}(t_{1},t_{2})$, defined by Eq. (17), and its conjugate $\hat{C}^{\phi}(t_{1},t_{2})$ to factorize the action across the neuron index $i$. This leads to a partition function for a statistical field theory involving $C^{\phi}(t_{1},t_{2})$ and $\hat{C}^{\phi}(t_{1},t_{2})$:
@@ F.1 Two-point functions
where $\left\langle\cdots\right\rangle_{\star}$ denotes an average within the dynamic process described by $W[C_{\star}^{\phi},0]$. This recovers the same single-site process [Eqs. (70) and (71)] and self-consistency condition [Eq. (72)] described in Appendix E.